# Parameter-Efficient Fine-Tuning in Large Models: A Survey of Methodologies

### v2

@@ II-C2 MLLM: Architecture
II-D Instruction Tuning
Instruction tuning in large language models has undergone significant development, evolving from initial efforts in multi-task fine-tuning without explicit instruction prompts to sophisticated techniques leveraging diverse tasks and templates. Early work focused on improving downstream task performance through large-scale multi-task fine-tuning [54, 55, 56, 57], while other efforts [58, 59, 60] converted a range of NLP tasks into a single generative question answering format using prompt instructions. The instruction tuning began in 2020 with the release of several task collections, including Natural Instructions [61], Flan 2021 [62], and PromptSource [63]. These collections aggregated large NLP datasets and provided templatized instructions for zero-shot prompting, enabling models to generalize to unseen instructions. MetaICL [64] emphasized few-shot prompting without explicit instructions, using input-output examples to teach tasks in-context. Research confirmed the benefits of task and template diversity, with some studies highlighting the advantages of inverting inputs and outputs to create new tasks [64]. The subsequent phase saw the expansion and combination of resources, with collections like SuperNatural Instructions [65] and OPT-IML [66] integrating more datasets and tasks. This phase also introduced multilingual instruction tuning, as seen in xP3 [67], and incorporated Chain-of-Thought training prompts in Flan 2022 [68]. These expanded collections included most tasks from previous resources, establishing a strong foundation for future open-source work. Current and future research is exploring new directions, such as synthetic data generation for creative and open-ended dialogue tasks [69, 70, 71, 72] and integrating human feedback on model responses [34, 73, 74, 41, 75]. These approaches are viewed as complementary to foundational instruction tuning methods, driving further advancements in the field.
@@ II-D Instruction Tuning
II-E Alignment Tuning and RLHF
@@
Figure 2: RLHF Workflow: This figure is from InstructGPT
@@
  • Supervised Fine-Tuning (SFT). Initially, a supervised dataset consisting of input prompts and desired outputs is applied to fine-tune the language model. These prompts and outputs can be written by human labelers for some specific tasks while ensuring the diversity of tasks. This step helps the model learn expected behaviors.
@@
  • Reward model training. A reward model is trained using human feedback data. The LLM is employed to generate a certain number of output texts using sampled prompts as input. Then human labelers rank these output pairs based on their preferences. Given human predictions, the reward model is trained to predict these rankings, effectively learning human preferences. Notably, [81] proposes an approach, namely Reinforcement Learning from AI Feedback (RLAIF), the annotation of preference on response pairs can be generated by an AI agent, increasing the automatic ability of the reinforcement process.
@@
  • Reinforcement learning fine-tuning: The final step involves formalizing the alignment process as a reinforcement learning problem. Here, the pre-trained language model acts as a policy generating text, with the reward model providing feedback scores. To prevent the model from deviating too far from its initial state, a penalty term is often included in the reward function. The language model is then optimized using algorithms like SARSA [82], DQN [83], PPO [84], iteratively improving its performance based on human-aligned rewards.
II-F Dataset for LLM training and tuning
@@
A critical component of the development and deployment of LLMs is the datasets used at various stages of their lifecycle, which significantly influence their capabilities and performance. In this section, we delve into the datasets that are instrumental in the pre-training, tuning, and other auxiliary phases of LLMs. The pre-training phase is where an LLM absorbs the foundational knowledge from a diverse array of textual data. This stage is pivotal, as it sets the stage for the model’s general understanding of language. The datasets used in pre-training are vast and varied, encompassing everything from the sprawling expanse of the internet to curated collections of literature and encyclopedias. Tuning is a process where the LLM is fine-tuned on specific tasks or domains. This phase refines the model’s abilities, enabling it to perform with greater precision and relevance in targeted applications. Tuning datasets are often more specialized and may include annotated examples that guide the model towards desired behaviors and outputs. Lastly, other datasets serve a multitude of purposes throughout the LLM’s development. These may include evaluation datasets that help assess the model’s performance, auxiliary datasets that broaden its knowledge base, or datasets that are used to imbue the model with specific skills or competencies.
@@
Commonly Used Datasets for Pre-training. In the realm of LLMs, the pre-training phase is instrumental in establishing a robust foundation upon which the model’s linguistic prowess is built. LLMs, with their exponentially larger parameter counts, necessitate an extensive and diverse corpus of training data that spans a multitude of topics and linguistic expressions. This data not only serves as the bedrock for the model’s comprehension of language but also influences its ability to generalize and adapt to new contexts and tasks. To meet these requirements, a variety of comprehensive and accessible datasets have been curated and made available for the research community. In this section, we embark on an overview of the datasets that are pivotal in the pre-training of LLMs. We categorize these datasets based on the type of content they provide, which can be broadly divided into seven distinct groups: Webpages, Books, Code, Social Media, Wikipedia, and a diverse array of other sources. Each of these categories contributes unique elements to the model’s knowledge base, ensuring a well-rounded understanding of human language and its myriad uses.
Collections Categories Publication Time Size
Common Crawl[85] Webpages 2023 400TB
WuDaoCorpora-Text[86] Webpages 2022 5TB
BookCorpusOpen[87] Books 2015 9.05GB
PG-19[88] Books 2020 11.74GB
BIGQUERY[89] Code 2023 341.1GB
The Stack[90] Code 2023 3TB
OpenWebText[91] Social Media 2019 38GB
Pushshift Reddit[92] Social Media 2020 89.1GB
wikipedia[93] wikipedia 2023 71.8GB
The Pile[94] Others 2020 800GB
S2ORC[95] Others 2020 80.5GB
MultiUN[96] Others 2010 31.8GB
TABLE II: A detailed list of available datasets for pre-training
Commonly Used Datasets for tuning. The fine-tuning phase of LLMs is a critical juncture where models are honed to perform with greater precision on specific tasks or domains. This process leverages a variety of datasets designed to target the skills and knowledge that the model needs to develop. In this table, we show some of the key datasets used for fine-tuning LLMs, focusing on their characteristics, the tasks they support, and their impact on the models’ performance.
Collections Categories Publication Time Examples
GLUE[97] NLP Task 2018 1,485,043
SuperGLUE[98] NLP Task 2019 196,309
E2E NLG Challenge[99] NLP Task 2020 50k
WikiSQL[100] NLP Task 2017 80,654
WebNLG[101] NLP Task 2017 27,731
SAMSum[102] Daily Chat 2019 16,369
OASST1[103] Daily Chat 2023 161,443
WMT[104] Others 2019 124,448,248
XSUM[105] Others 2018 200,000
DART[106] Others 2021 82,000
TABLE III: A detailed list of available datasets for tuning
Other Datasets. Beyond the foundational pretraining and targeted tuning phases, the realm of LLMs is enriched by a multitude of other datasets. These datasets, while not central to the initial training or fine-tuning processes, are indispensable for expanding the model’s expertise and enhancing its performance across a spectrum of nuanced language tasks. They embody the diversity of linguistic challenges and provide a testing ground for models to demonstrate adaptability and depth in understanding.In the following sections, we will explore these other datasets in detail, highlighting their unique characteristics, the specific tasks they target, and the value they bring to the comprehensive development of LLMs.
Collections Categories Publication Time Examples
HH-rlhf[74] Dialogue and Preference 2022 169,000
PKU-SafeRLHF[107] Dialogue and Preference 2023 362,000
HotpotQA[108] Question-Answering 2018 113,000
SHP[109] Community Preference 2022 385,000
MMLU[110] Knowledge Testing 2021 15,908
ARC[111] Knowledge Testing 2018 7,787
TABLE IV: Other datasets
@@ III PEFT Taxonomy
PEFT techniques are typically divided into three primary categories: Additive PEFT (III-A), which introduces additional trainable components or parameters into the pre-existing model; Reparameterized PEFT (III-B), a method that restructures the model’s parameters for the training phase and then reverts to the original form for inference; and Selective PEFT (III-C), which targets the optimization of a specific subset of the model’s parameters. The main differences between them are illustrated in Figure 3. Beyond these, Hybrid PEFT (III-D), which amalgamates the strengths of various PEFT approaches. In addition to these, there are specialized adaptations such as Quantization PEFT (III-E) tailored for the quantization process, and Multi-task PEFT (III-F) aimed at enhancing multi-task learning capabilities. A comprehensive classification of PEFT methods is depicted in Figure 4.
Figure 3: Illustration of different types of PEFT methods.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6.0em,font=, where level=2text width=5.5em,font=, where level=3text width=5.5em,font=, [ PEFT Methods for PLMs, ver [ Additive [ Adapter [ Sequential Adapter [15], Residual Adapter [112], AdapterDrop [113], Tiny-Attn Adapter [114],
Parallel Adapter [115], CIAT [116], CoDA [117], Hadamard Adapter [118], Compacter [119],
SparseAdapter [120] , leaf, text width=34.1em ] ] [ Soft Prompt [ Prefix-tuning [16], p-tuning [121], p-tuning v2 [122], prompt-tuning [17], SMoP [123], APT [124],
IDPG [125], LPT [126], SPT [127], APrompt [128], DePT [129], InfoPrompt [130], Xprompt [131],
@@
Decomposition [ Intrinsic SAID [139], LoRA [19], KronA [140] , leaf, text width=34.1em ] ] [ LoRA
Derivatives [ Dynamic Rank [ DyLoRA [141], AdaLoRA [20], IncreLoRA [142], SoRA [143] , leaf, text width=27em ] ] [ LoRA
Improvement [ LoRA+ [144], LoRA-FA [145], DoRA [146], Laplace-LoRA [147],
Kernel-mix [148], PeriodicLoRA [149], HydraLoRA [150], AFLoRA [151],
@@
Masking [ U-Diff pruning [154], U-Bitfit [155], PaFi [156], FishMask [157], Fish-Dip [158], LT-SFT [159],
SAM [160], Child-tuning [161], U-MAM [155], Threshold-Mask [162], LoRAPrune [163] , leaf, text width=34.1em ] ] [ Structural
Masking [ S-Diff pruning [154], S-Bitfit [155], FAR [164], Bitfit [22], Xattn Tuning [165], SPT [166],
@@
ProPETL [173] , leaf, text width=41.2em ] ] [ Quantization [ BI-Adapter [174], PEQA [175], QLoRA [176], LQ-LoRA [177], QA-LoRA [178], QDyLoRA [179], LoftQ [180],
BitDelta [181] , leaf, text width=41.2em ] ] [ Multi-task [ Adapter
-based [ AdapterFusion [182], AdaMix [183], PHA [184], AdapterSoup [185], MerA [186], Hyperformer [187] , leaf, text width=34.1em ] ] [ Soft Prompt
-based [ SPoT [188], ATTEMPT [189], MPT [190], IPT [191], TPT [192] , leaf, text width=34.1em ] ] [ LoRA
-based [ LoRAHub [193], MOELoRA [194], L-LoRA [195], MTLoRA [196] , leaf, text width=34.1em ] ] ] ]

### v3

@@ II-A3 Multimodal Large Language Models
II-B Optimization , Datasets, and Evaluation of Large Language Models
II-B1 Instruction Tuning
Instruction tuning in large language models has undergone significant development, evolving from initial efforts in multi-task fine-tuning without explicit instruction prompts to sophisticated techniques leveraging diverse tasks and templates. Early work focused on improving downstream task performance through large-scale multi-task fine-tuning [76, 77, 78, 79], while other efforts [80, 81, 82] converted a range of NLP tasks into a single generative question answering format using prompt instructions. The instruction tuning began in 2020 with the release of several task collections, including Natural Instructions [83], Flan 2021 [84], and PromptSource [85]. These collections aggregated large NLP datasets and provided templatized instructions for zero-shot prompting, enabling models to generalize to unseen instructions. MetaICL [86] emphasized few-shot prompting without explicit instructions, using input-output examples to teach tasks in-context. Research confirmed the benefits of task and template diversity, with some studies highlighting the advantages of inverting inputs and outputs to create new tasks [86]. The subsequent phase saw the expansion and combination of resources, with collections like SuperNatural Instructions [87] and OPT-IML [88] integrating more datasets and tasks. This phase also introduced multilingual instruction tuning, as seen in xP3 [89], and incorporated Chain-of-Thought training prompts in Flan 2022 [90]. These expanded collections included most tasks from previous resources, establishing a strong foundation for future open-source work. Current and future research is exploring new directions, such as synthetic data generation for creative and open-ended dialogue tasks [91, 92, 93, 94] and integrating human feedback on model responses [34, 95, 53, 96]. These approaches are viewed as complementary to foundational instruction tuning methods, driving further advancements in the field.
@@ II-B1 Instruction Tuning
II-B2 Alignment Tuning and RLHF
@@
Figure 2: RLHF Workflow: This figure is from InstructGPT, which interprets the RL process.
@@
  • Supervised Fine-Tuning (SFT): Initially, a supervised dataset consisting of input prompts and desired outputs is applied to fine-tune the language model. These prompts and outputs can be written by human labelers for some specific tasks while ensuring the diversity of tasks. This step helps the model learn expected behaviors.
@@
  • Reward Model Training: A reward model is trained using human feedback data. The LLM is employed to generate a certain number of output texts using sampled prompts as input. Then human labelers rank these output pairs based on their preferences. Given human predictions, the reward model is trained to predict these rankings, effectively learning human preferences. Notably, [102] proposes an approach, namely Reinforcement Learning from AI Feedback (RLAIF), the annotation of preference on response pairs can be generated by an AI agent, increasing the automatic ability of the reinforcement process.
@@
  • Reinforcement Learning Fine-Tuning: The final step involves formalizing the alignment process as a reinforcement learning problem. Here, the pre-trained language model acts as a policy generating text, with the reward model providing feedback scores. To prevent the model from deviating too far from its initial state, a penalty term is often included in the reward function. The language model is then optimized using algorithms like SARSA [103], DQN [104], PPO [105], DPO [106], and GRPO [63], iteratively improving its performance based on human-aligned rewards.
II-B3 Datasets for LLM
@@
A critical component of the development and deployment of LLM is the datasets used at various stages of their lifecycle, which significantly influence their capabilities and performance. In this section, we delve into the datasets that are instrumental in the Pre-training, SFT, and RLHF. The Pre-training phase is where an LLM absorbs the foundational knowledge from a diverse array of textual data. This stage is pivotal, as it sets the stage for the model’s general understanding of language. The datasets used in Pre-training are vast and varied, encompassing everything from the sprawling expanse of the internet to curated collections of literature and encyclopedias. SFT is the process where the LLM is fine-tuned on specific tasks or domains. This phase refines the model’s abilities, enabling it to perform with greater precision and relevance in targeted applications. SFT datasets are often more specialized and may include annotated examples that guide the model towards desired behaviors and outputs. RLHF is the stage where the LLM is further optimized based on human feedback. This phase enhances the model’s alignment with human preferences and values, ensuring that its outputs are more aligned with user expectations. RLHF datasets typically consist of human-labeled examples and feedback, which help the model learn to prioritize high-quality and contextually appropriate responses.
@@
Commonly Used Datasets for Pre-training. In the realm of LLM, the pre-training phase is instrumental in establishing a robust foundation upon which the model’s linguistic prowess is built. LLM, with their exponentially larger parameter counts, necessitate an extensive and diverse corpus of training data that spans a multitude of topics and linguistic expressions. This data not only serves as the bedrock for the model’s comprehension of language but also influences its ability to generalize and adapt to new contexts and tasks. To meet these requirements, a variety of comprehensive and accessible datasets have been curated and made available for the research community.
In this section, we embark on an overview of the datasets that are pivotal in the pre-training of LLM. We categorize these datasets based on the type of content they provide, which can be broadly divided into seven distinct groups: Webpages, Books, Code, Social Media, Wikipedia, and a diverse array of other sources. Each of these categories contributes unique elements to the model’s knowledge base, ensuring a well-rounded understanding of human language and its myriad uses. Here are 2 typical Pre-training Datasets and their importance in evaluating PEFT Methods:
  • Common Crawl: The Common Crawl corpus is an extensive, unstructured, multilingual dataset of webpages, encompassing over eight years of web crawler data. This dataset is available in various formats, including web archive, web archive transformation, and web-extracted text. Many pre-training corpora are obtained through data preprocessing based on this corpus, which provides a vast and diverse source of text for language models. Its unstructured nature and multilingual content make it an ideal resource for training models that need to handle a wide variety of text types and languages. Importantly, the Common Crawl corpus plays a crucial role in evaluating PEFT methods. Its vast and varied content provides a comprehensive base for pre-training models that can then be fine-tuned using PEFT techniques. This allows researchers to assess how effectively PEFT methods can enhance model performance across diverse linguistic contexts.
  • The Pile: The Pile is a large-scale, diverse language modeling dataset consisting of 22 data subsets, designed to capture text in as many forms as possible and cover a wide range of textual content. The corpus includes academic papers, code, legal materials, patents, subtitles, chat content, parallel corpora, and more. This diversity ensures that models trained on The Pile are exposed to a broad spectrum of language use cases, making them more adaptable to various downstream tasks. In the context of evaluating PEFT methods, The Pile offers a robust testbed. Its rich diversity of text types allows researchers to evaluate how well these fine-tuning methods can adapt models to different domains and tasks, thereby enhancing their understanding of the effectiveness of PEFT methods in various applications.
Collections Categories Publication Time Size URL
Common Crawl [^4] https://commoncrawl.org/ Webpages 2023 400TB https://commoncrawl.org/
WuDaoCorpora-Text [107] Webpages 2023 5TB https://data.baai.ac.cn/details/WuDaoCorporaText
BookCorpusOpen  [108] Books 2015 9.05GB https://huggingface.co/datasets/defunct-datasets/bookcorpusopen
PG-19  [109] Books 2020 11.74GB https://huggingface.co/datasets/deepmind/pg19
The Stack [110] Code 2022 3TB https://huggingface.co/datasets/bigcode/the-stack
OpenWebText [111] Social Media 2019 38GB https://huggingface.co/datasets/Skylion007/openwebtext
Pushshift Reddit [112] Social Media 2020 89.1GB https://zenodo.org/records/3608135
Wikipedia [^5] https://www.wikipedia.org/ Wikipedia 2023 71.8GB https://huggingface.co/datasets/wikimedia/wikipedia
The Pile [113] Others 2020 800GB https://pile.eleuther.ai/
S2ORC [114] Others 2020 80.5GB https://huggingface.co/datasets/sentence-transformers/s2orc
MultiUN [115] Others 2010 31.8GB https://huggingface.co/datasets/Helsinki-NLP/multiun
TABLE II: A Curated List of Datasets for Pre-Training. This table provides a comprehensive overview of various datasets used for pre-training purposes in natural language processing tasks. It includes details such as the collection name, the corpus it belongs to, publication year, size in terms of tokens, and the URL for accessing the dataset. The datasets listed cover a range of sources from web pages to books, offering a diverse set of data for training models in different domains.
Commonly Used Datasets for SFT and RLHF. Two critical stages in LLM are SFT and RLHF. These stages are designed to enhance the model’s performance on specific tasks and align its outputs with human preferences. This section provides an overview of these two stages, highlighting their significance and the datasets used to support them.
SFT is a process where LLM are trained on specialized datasets to improve their performance on specific tasks. This stage is crucial for adapting the model to particular domains or applications. SFT involves using annotated datasets that provide examples of desired outputs for given inputs. By training on these datasets, the model learns to generate more accurate and contextually relevant responses. RLHF is particularly effective in enhancing the model’s ability to follow human instructions. These datasets provide a comprehensive set of examples that help the model learn to discern correct answers from plausible alternatives.
Collections Categories Publication Time Examples URL
E2E NLG [116] NLP Task 2020 50,000 https://sites.google.com/site/hwinteractionlab/E2E/
WikiSQL [117] NLP Task 2017 80,654 https://huggingface.co/datasets/Salesforce/wikisql
WebNLG [118] NLP Task 2017 27,731 https://huggingface.co/datasets/web_nlg
SAMSum [119] Daily Chat 2019 16,369 https://huggingface.co/datasets/Samsung/samsum
OASST1 [120] Daily Chat 2023 161,443 https://huggingface.co/datasets/OpenAssistant/oasst1
WMT [^6] https://www.statmt.org/wmt19/ Others 2019 124,448,248 https://huggingface.co/datasets/wmt/wmt19
XSUM [121] Others 2018 200,000 https://huggingface.co/datasets/EdinburghNLP/xsum
DART [122] Text Generation 2021 82,000 https://github.com/Yale-LILY/dart
HH-rlhf [123] Dialogue and Preference 2022 169,000 https://huggingface.co/datasets/Anthropic/hh-rlhf
PKU-SafeRLHF [124] Dialogue and Preference 2023 362,000 https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF
HotpotQA [125] Question-Answering 2018 113,000 https://huggingface.co/datasets/hotpotqa/hotpot_qa
SHP [126] Community Preference 2022 385,000 https://huggingface.co/datasets/stanfordnlp/SHP
TABLE III: A Curated List of datasets for SFT and RLHF. This table provides an overview of the datasets used in SFT and RLHF phases, categorized by their primary purposes and characteristics. This categorization helps in understanding the diversity and scope of data used to train and fine-tune models in different phases of development. The URLs provided allow researchers and practitioners to access these datasets for further analysis and experimentation.
II-B4 LLM evaluation
The burgeoning field of LLM research has necessitated the development of robust evaluation frameworks to accurately gauge the capabilities and limitations of these sophisticated AI systems. Evaluation serves multiple critical functions: it benchmarks model performance across a spectrum of tasks, identifies areas for improvement, and ensures that advancements in LLM technology align with ethical and practical standards. In the academic and professional realms of LLM evaluation, it is widely recognized that a multifaceted approach is essential to gauge the capabilities and limitations of these advanced AI systems comprehensively. The Qwen blog’s evaluation of the Qwen2.5 base language model [^7] https://qwenlm.github.io/blog/qwen2.5-llm/, underscore the importance of using multiple benchmarks to assess the model’s performance across various domains thoroughly.
Platforms such as Hugging Face offer a suite of datasets for this purpose [^8] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard, including IFEval, BBH, MATH [127], GPQA [128], and MUSR  [129]. These datasets encompass a broad spectrum of tasks, ranging from language modeling to problem-solving in mathematics, ensuring a comprehensive evaluation of model competencies. Models like Qwen2.5 is evaluated using a diverse array of datasets that cover general tasks such as MMLU, and HellaSwag, as well as specialized tasks in math and science with datasets like GPQA and MATH, and coding tasks including HumanEval and MBPP. Additionally, multilingual capabilities are assessed through datasets like Multi-Exam and Multi-Translation.
To achieve a comprehensive evaluation of a language model’s performance, it is often necessary to employ a combination of benchmarks. These benchmarks should be representative of real-world scenarios and cover diverse domains and linguistic complexities. The evaluations include a variety of tests that measure the model’s ability to handle extended dialogues and manage a variety of tasks. By leveraging these diverse datasets and assessments, researchers can effectively benchmark LLM and guide their development towards practical applications, ensuring alignment with ethical and practical standards. This section explores benchmarking in two parts: general tasks and specialized tasks.
  • General Tasks: General tasks are designed to assess the broad capabilities of LLM across a wide range of subjects and skills. These benchmarks are essential for evaluating the foundational knowledge and general reasoning abilities of LLM. These benchmarks help determine how well models can understand and generate text in various contexts, ensuring that they possess a solid understanding of language fundamentals. Datasets such as MMLU, ARC, and HellaSwag are commonly used for general evaluations.
  • Specialized Tasks: Specialized tasks focus on evaluating LLM in specific domains, such as mathematics, coding, and natural language understanding. These benchmarks are designed to assess the model’s proficiency in particular areas, providing a deeper understanding of their specialized skills. Specialized tasks are crucial for identifying domain-specific strengths and weaknesses, ensuring that models can effectively apply their knowledge in practical scenarios.
Collections task Publication Time examples URL
MMLU [130] general 2021 15,908 https://huggingface.co/datasets/cais/mmlu
ARC [131] general 2018 7,787 https://huggingface.co/datasets/allenai/ai2_arc
HellaSwag [132] general 2019 59,950 https://huggingface.co/datasets/Rowan/hellaswag
GLUE [133] natural language understanding 2018 1,485,043 https://huggingface.co/datasets/nyu-mll/glue
SuperGLUE [134] natural language understanding 2019 196,309 https://huggingface.co/datasets/aps/super_glue
GSM8K [135] Science and mathematics 2021 17,584 https://huggingface.co/datasets/openai/gsm8k
Theoremqa [136] Science and mathematics 2023 800 https://huggingface.co/datasets/TIGER-Lab/TheoremQA
Humaneval [52] code 2021 164 https://huggingface.co/datasets/openai/openai_humaneval
MBPP [137] code 2021 1,401 https://huggingface.co/datasets/google-research-datasets/mbpp
AGIEval [138] Exam 2023 8,062 https://github.com/ruixiangcui/AGIEval
GAOKAO-Bench [139] Exam 2023 2,811 https://github.com/OpenLMLab/GAOKAO-Bench
TruthfulQA [47] other 2021 1,634 https://huggingface.co/datasets/truthfulqa/truthful_qa
BBH [140] other 2022 6,511 https://huggingface.co/datasets/lukaemon/bbh
TABLE IV: A typical list of available datasets for LLM Evaluation. This table provides an exhaustive compilation of datasets pertinent to the evaluation of LLM. These datasets span a diverse array of tasks, from general to domain-specific, aiming to holistically assess the performance of LLM across various scenarios. The table delineates the publication timeline, the number of examples, and the access points (URLs) for each dataset, facilitating researchers in procuring and utilizing these resources.
@@ III PEFT Taxonomy
PEFT techniques are typically divided into three primary categories: Additive PEFT (III-A), which introduces additional trainable components or parameters into the pre-existing model; Reparameterized PEFT (III-B), a method that restructures the model’s parameters during the training phase and then reverts to the original form for inference; and Selective PEFT (III-C), which focuses on optimizing a specific subset of the model’s parameters. Besides these, there is the Hybrid PEFT (III-D), which combines the strengths of various PEFT approaches. Additionally, there are specialized adaptations such as Quantization PEFT (III-E) designed for the quantization process, and Multi-task PEFT (III-F) aimed at enhancing multi-task learning capabilities. A conceptual illustration of the core principles underlying these PEFT methodologies is presented in Figure 3. A comprehensive classification of PEFT methods is depicted in Figure 4. The main ideas, number of trainable parameters, applications, and limitations of different types of PEFT methods are summarized in Table V. To facilitate a more intuitive understanding of the performance differences among various PEFT methods, Table VI presents the performance results of representative PEFT methods of different types across various base models and tasks.
Figure 3: Illustration of the main idea of different types of PEFT methods
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=, where level=2text width=7.0em,font=, where level=3text width=7.0em,font=, [ PEFT Methods for PLMs, ver [ Additive [ Adapter [ Sequential Adapter [15], Residual Adapter [141],
AdapterDrop [142], Tiny-Attn Adapter [143],
Parallel Adapter [144], CIAT [145], CoDA [146],
Hadamard Adapter [147], Compacter [148],
SparseAdapter [149] , leaf, text width=35.6em ] ] [ Soft Prompt [ Prefix-tuning [16], p-tuning [150],
p-tuning v2 [151], prompt-tuning [17],
SMoP [152], APT [153], IDPG [154],
LPT [155], SPT [156], APrompt [157],
DePT [158], InfoPrompt [159], Xprompt [160],
@@
PASTA [164] , leaf, text width=35.6em ] ] [ Others [ IPA [165], LST [166], Attention-Fusion [167] , leaf, text width=35.6em ] ] ] [ Reparameterized [ Low-rank
Decomposition [ Intrinsic SAID [168], LoRA [19],
KronA [169] , leaf, text width=35.6em ] ] [ LoRA
Derivatives [ Dynamic Rank [ DyLoRA [170], AdaLoRA [20],
IncreLoRA [171], SoRA [172] , leaf, text width=27em ] ] [ LoRA
Improvement [ LoRA+ [173], LoRA-FA [174],
DoRA [175], Laplace-LoRA [176],
Kernel-mix [177], PeriodicLoRA [178],
HydraLoRA [179], AFLoRA [180],
@@
Masking [ U-Diff pruning [183], U-Bitfit [184], PaFi [185],
FishMask [186], Fish-Dip [187], LT-SFT [188],
SAM [189], Child-tuning [190], U-MAM [184],
Threshold-Mask [191], LoRAPrune [192] , leaf, text width=35.6em ] ] [ Structural
Masking [ S-Diff pruning [183], S-Bitfit [184],
FAR [193], Bitfit [22], Xattn Tuning [194],
SPT [195], S-MAM [184] , leaf, text width=35.6em ] ] ] [ Hybrid [ UniPELT [196], S4 [197], MAM Adapter [144],
LLM-Adapters [198], NOAH [199], AUTOPEFT [200],
@@
LQ-LoRA [206], QA-LoRA [207], QDyLoRA [208],
LoftQ [209], BitDelta [210] , leaf, text width=44.2em ] ] [ Multi-task [ Adapter
-based [ AdapterFusion [211], AdaMix [212],
PHA [213], AdapterSoup [214],
MerA [215], Hyperformer [216] , leaf, text width=35.6em ] ] [ Soft Prompt
-based [ SPoT [217], ATTEMPT [218], MPT [219],
IPT [220], TPT [221] , leaf, text width=35.6em ] ] [ LoRA
-based [ LoRAHub [222], MOELoRA [223],
L-LoRA [224], MTLoRA [225] , leaf, text width=35.6em ] ] ] ]